OpenStudy (anonymous):

if f(x)=4x-7 and g(x)=2x+4, evaluate f(x)+g(x)for x=-3 a.-52 b.-21 c.-22 d.-15

OpenStudy (hoblos):

f(x)+g(x) = 4x-7 + 2x+4 = 6x -3 for x=-3 .... 6x -3= 6(-3)-3 = -18-3=-21